OpenStudy (anonymous):

5(x − 1) = 3(x + 2) − 7

OpenStudy (anonymous):

distribute the 5 into x-1 and the 3 to x-2 and combine the like terms

OpenStudy (diyadiya):

Distribute 5(x-1)=3(x+2)-7 5x-5=3x+6-7 Subtract 3x from both sides 2x-5=-1 Add 5 on both sides 2x=4 divide by 2 x=2